OpenStudy (anonymous):

The scale on a map indicates that 2 cm is equivalent to 15 miles. The distance Christine needs to travel is 5 cm on the map. How many miles does Christine need to travel? a) 3 b) 15 c) 30 d) 37.5

OpenStudy (ghazi):

2cm= 15 miles 1cm= 15/2 miles therefore 5cm= 15/2* 5 miles

OpenStudy (ghazi):

7.5*5....=

OpenStudy (anonymous):

d?